OpenStudy (girlgoyleh.):

What is the solution to the system? line 1 line 2 x y x y -2 -4 2 3 -1 -1 3 1

OpenStudy (dtan5457):

You have 2 points per line. Get the slope of each one, and plug them each into point slope formula y-y1=m(x-x1)

OpenStudy (girlgoyleh.):

for the first one, its -1-(-4) -1-(-2)

OpenStudy (girlgoyleh.):

which equals 3

OpenStudy (dtan5457):

So like for the first one (-2,-4)(-1,-1)= -3/-1= slope of 3 y+4=3(x+2) y+4=3x+6 y=3x+2

OpenStudy (dtan5457):

Can you get the line for the 2nd equation?

OpenStudy (girlgoyleh.):

-2

OpenStudy (girlgoyleh.):

so its 3,-2

OpenStudy (dtan5457):

what did you get for the line ?

OpenStudy (girlgoyleh.):

for line number 2, i got the answer -2

OpenStudy (dtan5457):

The slope is indeed -2. Now plug that into y-y1=m(x-x1)

OpenStudy (girlgoyleh.):

how do i do that?

OpenStudy (dtan5457):

Take a point like 2,3 from your chart y-3=-2(x-2) can you simplify that?

OpenStudy (girlgoyleh.):

not really..

OpenStudy (dtan5457):

y-3=-2x+4 add the 3 y=-2x+7

OpenStudy (dtan5457):

Now you have two lines y=3x+2 y=-2x+7 you can solve for x by equating these together 3x+2=-2x+7 5x=5 x=1 plug that in any equation y=3x+2 y=3(1)+2 y=5 your solution=1,5

OpenStudy (girlgoyleh.):

wow youre good

OpenStudy (dtan5457):

thank you :)

OpenStudy (girlgoyleh.):

no thank you! hey, if i need help with a couple more questions, would you be willing?

OpenStudy (dtan5457):

ask it on a new question and i'll try, im tight on time though